1. If the sum of the first 20 terms of an A.P is 200, and the first term is 40, find the common difference.

2. Find the sum to infinity of a G.P whose first term is 2 and common ratio is 19/20

3. Find the sum of the first 10 terms of a G.P whose first term is 5 and the common ratio is 2
1
Expert's answer
2019-05-29T09:09:33-0400

1) "S_n=\frac{n(2a+(n-1)d)n}{2}. \;\;S_{20}=\frac{(80+19d)*20}{2}=200."

So, "d=\frac{20-80}{19}=-\frac{60}{19}."


2) "S=\frac{a}{1-r}=\frac{2}{1-\frac{19}{20}}=40."


3) "S_n=\frac{a(1-r^n)}{1-r}."

"S_{10}=\frac{5(1-2^{10})}{1-2}=5115."